DIRC Operations Weekly Report

Period: Wednesday 2007/7/18 - Wednesday 2007/7/25


Oncall: Rolf Andreassen
Fast Monitoring: Rolf Andreassen, Justine Serrano

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: ~1.0 x 1034 (last week: ~9.7 x 1033)

PEP delivery: 2.97 fb-1 (last weeks: 3.19 fb-1)

BaBar integrated lumi:2.81 fb-1 (last weeks: 3.12 fb-1)

Run 6 total delivery: 64.45 fb-1

Run 6 total BaBar recorded: 61.95 fb-1 total, including 4.18 fb-1 off-peak data

PEP activities/issues this week

  • Throughout : Problems with beam-current monitor led to difficulties in trickling the HER. This made it necessary to go back to the old fill-and-coast system for long periods, particularly almost the whole weekend.
  • 07/20: Replacement of QD4R rectifier took roughly 1.5 hours. Tuner in HER 12-4 stuck, 1.3 hours lost.
  • 07/22: Again the HER 12-4 tuner, 25 minutes to do 'local cable reseat'. Kicking the tuner until it unsticks?
  • 07/23: Swap chassis on EP02 magnet controller, 3 hours.
  • 07/24: Fix loose connection in region 6, which was causing RF aborts. Replace 12-4 cavity tuner driver's supply.

    BaBar activities/issues this week

  • 07/21: Maximum length of runs changed to 75 minutes to accommodate PEP fill-and-coast cycle.

    DIRC activities/issues this week

  • 07/19 : Matt turned off dump valve for the UPS battery replacement, but the wrong kind of batteries were delivered.
  • 07/24 : Matt cleaned the air filters.

    See DIRC logbook for details

    Lumi and Background this week

    Scheduled activities

    Pending Work

    Manpower situation

    Justine, Jose and Rolf rotating backup & on-call commissioning duties.

    Links